OpenStudy (anonymous):

Factor using difference of squares. x^15-64 Please explain Im just not sure what the cube root of x^15 is....

ganeshie8 (ganeshie8):

hint : \(\large a^{m\times n} = (a^m)^n\)

OpenStudy (anonymous):

sttill confused I'm also stuck on 24x^6-1029y^3

ganeshie8 (ganeshie8):

\(\large x^{15}-64 = x^{5\times 3} - 4^3 = (x^5)^3 - 4^3\)

ganeshie8 (ganeshie8):

You need to use difference of CUBES...

OpenStudy (anonymous):

yes thats what I meant.....

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

http://www.purplemath.com/modules/specfact2.htm